# TIP-I2V: A Million-Scale Real Text and Image Prompt Dataset for Image-to-Video Generation

### v1

@@ Abstract
Video generation models are revolutionizing content creation, with image-to-video models drawing increasing attention due to their enhanced controllability, visual consistency, and practical applications. However, despite their popularity, these models rely on user-provided text and image prompts, and there is currently no dedicated dataset for studying these prompts. In this paper, we introduce TIP-I2V, the first large-scale dataset of over $1.70$ million unique user-provided Text and Image Prompts specifically for Image-to-Video generation. Additionally, we provide the corresponding generated videos from five state-of-the-art image-to-video models. We begin by outlining the time-consuming and costly process of curating this large-scale dataset. Next, we compare TIP-I2V to two popular prompt datasets, VidProM (text-to-video) and DiffusionDB (text-to-image), highlighting differences in both basic and semantic information. This dataset enables advancements in image-to-video research. For instance, to develop better models, researchers can use the prompts in TIP-I2V to analyze user preferences and evaluate the multi-dimensional performance of their trained models; and to enhance model safety, they may focus on addressing the misinformation issue caused by image-to-video models. The new research inspired by TIP-I2V and the differences with existing datasets emphasize the importance of a specialized image-to-video prompt dataset. The project is publicly available at https://tip-i2v.github.io.
Figure 1: TIP-I2V is the first dataset comprising over 1.70 million unique user-provided text and image prompts. Besides the prompts, TIP-I2V also includes videos generated by five state-of-the-art image-to-video models ($\mathtt{Pika}$ [5], $\mathtt{Stable}$ $\mathtt{Video}$ $\mathtt{Diffusion}$ [8], $\mathtt{Open\text{-}Sora}$ [73], $\mathtt{I2VGen\text{-}XL}$ [71], and $\mathtt{CogVideoX\text{-}5B}$ [69]). The TIP-I2V contributes to the development of better and safer image-to-video models.
@@ 1 Introduction
The first Text and Image Prompt dataset for Image-to-Video generation (TIP-I2V). As shown in Fig. 1, our TIP-I2V dataset includes over $1.70$ million unique user-provided text and image prompts for image-to-video diffusion models, along with the corresponding generated videos, sourced from Pika Discord channels [5]. It is important to note that: (1) We intend to include videos generated by other state-of-the-art image-to-video diffusion models, including $\mathtt{Stable}$ $\mathtt{Video}$ $\mathtt{Diffusion}$ [8], $\mathtt{Open\text{-}Sora}$ [73], $\mathtt{I2VGen\text{-}XL}$ [71], and $\mathtt{CogVideoX\text{-}5B}$ [69]; however, due to limited computing resources, we only use $100,000$ randomly selected prompts to generate videos for each image-to-video model. Researchers are free to extend our TIP-I2V by generating more videos with these methods (or other state-of-the-arts, such as $\mathtt{Open\text{-}Sora\text{-}Plan}$ [31]) and our prompts; (2) we acknowledge that the currently generated videos are not perfect, and in the future, researchers are encouraged to use newly released image-to-video models (such as $\mathtt{Sora}$ [4] and $\mathtt{Movie}$ $\mathtt{Gen}$ [3]) and our prompts to further extend our TIP-I2V. Besides prompts and generated videos, our TIP-I2V also includes Universally Unique Identifier (UUIDs), timestamps, embeddings, subjects, and not-safe-for-work (NSFW) scores for these data points.
@@
Differences between TIP-I2V and other similar datasets in basic and semantic information. We notice that there are two popular prompt-gallery datasets in the visual generation community, i.e., VidProM [58] for text-to-video and DiffusionDB [64] for text-to-image. Our TIP-I2V mainly differs from them in: (1) Basic information: Both VidProM and DiffusionDB begin the generation with a text, while our TIP-I2V starts with a text and an image. (2) Semantics: Each text in our TIP-I2V focuses on how to bring static elements in the corresponding image to life through motion. For instance, some text prompts are “the hair and body movement, dancing”, “the flowers falling, the woman move”, and “make the statue broke down”. In contrast, the prompts in VidProM and DiffusionDB are more descriptive, i.e., they directly specify the content to be generated without referencing a specific object, such as “a dragon flying over a medieval city”, “a serene beach at sunset”, and “an astronaut walking on Mars”. The differences in basic and semantic information highlight a need for a specialized image-to-video prompt dataset.
@@
Exciting new research areas inspired by TIP-I2V. Our TIP-I2V helps researchers develop better and safer image-to-video diffusion models. For better models: (1) Enhancing user experience. Before TIP-I2V, researchers do not know which subjects users prefer to transform into videos and what directions they expect. However, with our dataset, researchers can cater to more users in the real world with limited training sources. This avoids the problem researchers previously faced, where blindly expanding the training set led to wasted resources and low user satisfaction. (2) Improving evaluation practicality. Existing image-to-video benchmarks, such as $\mathtt{VBench}$-$\mathtt{I2V}$ [25], $\mathtt{I2V}$-$\mathtt{Bench}$ [49], and $\mathtt{AIGCBench}$ [18], suffer from a limited number of topics and prompts designed by experts, which may not accurately reflect real-world user needs. With the help of TIP-I2V, researchers can build a more comprehensive and practical benchmark for evaluating image-to-video models. For safer models: A major safety concern of image-to-video generation is misinformation, i.e., they can make objects or humans in images to perform actions they never did. For instance, given an image of $\mathtt{Elon}$ $\mathtt{Musk}$ and $\mathtt{Donald}$ $\mathtt{Trump}$ together, image-to-video models could generate a video of them fighting, misleading the public. To address this issue, our TIP-I2V allows researchers to train a model to (1) distinguish between generated videos from images and real videos, and (2) trace the source image from any given frame in a generated video. Beyond these areas, we also encourage researchers to explore additional directions inspired by the TIP-I2V.
@@
Figure 2: A data point in our TIP-I2V includes UUID, timestamp, text and image prompt, subject, NSFW status of text and image, text and image embedding, and the corresponding generated videos.
@@ 3 Curating TIP-I2V
Fig. 2 is an example data point in the proposed TIP-I2V. It includes a UUID, timestamp, text and image prompts, subject, NSFW status for text and images, embeddings for both text and images, as well as generated videos from five state-of-the-art image-to-video diffusion models. The following explains how we collect these pieces of information.
@@
Assigning UUIDs, timestamps, subjects, embeddings and NSFW status. To facilitate subsequent research utilizing our TIP-I2V, we (1) calculate UUIDs based on unique prompts, (2) extract timestamps from Pika videos, (3) infer subjects using $\mathtt{GPT}$-$\mathtt{4o}$ [42], (4) embed text and image prompts using $\mathtt{text}$-$\mathtt{embedding}$-$\mathtt{3}$-$\mathtt{large}$ [41] and $\mathtt{CLIP}$ [47], respectively, and (5) assign NSFW status to text and image prompts using $\mathtt{Detoxify}$ [54] and $\mathtt{nsfw\_image\_detection}$ [17], respectively.
@@ 5.5 Other Promising Research
$\bullet$ Retrieval-augmented generation. Enhancing image-to-video models with retrieval can improve generation quality and relevance. By retrieving similar past prompts or generated videos in TIP-I2V, models may use this additional context to produce more accurate and coherent videos.
@@ 7 Comparing TIP-I2V with Panda-70M
As shown in Fig. 8, we compare the text semantics of our TIP-I2V and Panda-70M [12] using WizMap to highlight their differences.
@@ 14 Details of Fine-tuning VideoMAE
We fine-tune the Video Masked Autoencoder (VideoMAE) [53] on the TIP-ID dataset for a video classification task. Specifically, our preprocessing pipeline includes (1) temporal subsampling, (2) spatial transformations, and (3) normalization. During training, the spatial transformations consist of random short-side scaling, random cropping to $224\times 224$, and random horizontal flipping; for testing, we only resize frames to $224\times 224$. The pre-trained model is downloaded from https://huggingface.co/MCG-NJU/videomae-base, and we adjust it to match the number of classes in our dataset, i.e., 3, by updating the classification head. Training is distributed across a server with $8$ A100 GPUs. The mini-batch size is $8$, the learning rate is $5\times 10^{-5}$, and the number of iterations is $20,000$.
@@ 17 Potential Social Impact
The TIP-I2V dataset has potential for positive social impact by enhancing digital creativity and fostering responsible AI use. Specifically, by helping the creation of more advanced and user-responsive image-to-video models, TIP-I2V enables artists, content creators, and educators to create engaging and customized videos. Additionally, TIP-I2V contributes to the development of detection models that help verify authenticity, trace image sources, and prevent harmful content misuse. Nevertheless, the TIP-I2V dataset may also have potential negative social impacts if misused. Below, we outline several potential negative social impacts and provide corresponding solutions:
@@
$\bullet$ Privacy. Although, per Pika’s regulations, users agree to make their input and output publicly available, some may still feel uncomfortable with the inclusion of their prompts in TIP-I2V. To enhance user privacy, we implement the following measures: (1) each prompt is assigned a new UUID instead of user-identifiable original IDs; and (2) users have the right to request that their contributions be removed from TIP-I2V. They can simply email us to make this request.

### v2

@@ Abstract
Video generation models are revolutionizing content creation, with image-to-video models drawing increasing attention due to their enhanced controllability, visual consistency, and practical applications. However, despite their popularity, these models rely on user-provided text and image prompts, and there is currently no dedicated dataset for studying these prompts. In this paper, we introduce TIP-I2V, the first large-scale dataset of over $1.70$ million unique user-provided Text and Image Prompts specifically for Image-to-Video generation. Additionally, we provide the corresponding generated videos from five state-of-the-art image-to-video models. We begin by outlining the time-consuming and costly process of curating this large-scale dataset. Next, we compare TIP-I2V to two popular prompt datasets, VidProM (text-to-video) and DiffusionDB (text-to-image), highlighting differences in both basic and semantic information. This dataset enables advancements in image-to-video research. For instance, to develop better models, researchers can use the prompts in TIP-I2V to analyze user preferences and evaluate the multi-dimensional performance of trained models; and to enhance model safety, they may focus on addressing the misinformation issue caused by image-to-video models. The new research inspired by TIP-I2V and the differences with existing datasets emphasize the importance of a specialized image-to-video prompt dataset. The project is available at https://tip-i2v.github.io/.
Figure 1: TIP-I2V is the first dataset comprising over 1.70 million unique user-provided text and image prompts. Besides the prompts, TIP-I2V also includes videos generated by five state-of-the-art image-to-video models ($\mathtt{Pika}$ [5], $\mathtt{Stable}$ $\mathtt{Video}$ $\mathtt{Diffusion}$ [8], $\mathtt{Open\text{-}Sora}$ [73], $\mathtt{I2VGen\text{-}XL}$ [71], and $\mathtt{CogVideoX\text{-}5B}$ [69]). The TIP-I2V inspires new research directions related to image-to-video generation.
@@ 1 Introduction
The first Text and Image Prompt dataset for Image-to-Video generation (TIP-I2V). As shown in Fig. 1, our TIP-I2V dataset includes over $1.70$ million unique user-provided text and image prompts for image-to-video diffusion models, along with the corresponding generated videos, sourced from Pika Discord channels [5]. It is important to note that: (1) We intend to include videos generated by other state-of-the-art image-to-video diffusion models, including $\mathtt{Stable}$ $\mathtt{Video}$ $\mathtt{Diffusion}$ [8], $\mathtt{Open\text{-}Sora}$ [73], $\mathtt{I2VGen\text{-}XL}$ [71], and $\mathtt{CogVideoX\text{-}5B}$ [69]; however, due to limited computing resources, we only use $100,000$ randomly selected prompts to generate videos for each image-to-video model. Researchers are free to extend our TIP-I2V by generating more videos with these methods (or other state-of-the-arts, such as $\mathtt{Open\text{-}Sora\text{-}Plan}$ [31]) and our prompts; (2) we acknowledge that the currently generated videos are not perfect, and in the future, researchers are encouraged to use newly released image-to-video models (such as $\mathtt{Sora}$ [4]) and our prompts to further extend our TIP-I2V. Besides prompts and generated videos, our TIP-I2V also includes Universally Unique Identifier (UUIDs), anonymous UserIDs, timestamps, embeddings, subjects, and not-safe-for-work (NSFW) scores for these data points.
@@
Differences between TIP-I2V and other similar datasets in basic and semantic information. We notice that there are two popular prompt-gallery datasets in the visual generation community, i.e., VidProM [58] for text-to-video and DiffusionDB [64] for text-to-image. Our TIP-I2V mainly differs from them in: (1) Basic information: Both VidProM and DiffusionDB begin the generation with a text, while our TIP-I2V starts with a text and an image. (2) Semantics: Each text in our TIP-I2V focuses on how to bring static elements in the corresponding image to life through motion. For instance, some text prompts are “the hair and body dance”, “the flowers fall, the woman move”, and “make the statue break down”. In contrast, the prompts in VidProM and DiffusionDB are more descriptive, i.e., they directly specify the content to be generated without referencing a specific object, such as “a dragon flies over a city”, “a beach at sunset”, and “an astronaut walks on Mars”. The differences in basic and semantic information highlight a need for an image-to-video prompt dataset.
@@
Exciting new research areas inspired by TIP-I2V. Our TIP-I2V helps researchers develop better and safer image-to-video diffusion models. For better models: (1) Enhancing user experience. Before TIP-I2V, researchers do not know which subjects users prefer to transform into videos and what directions they expect. However, after analyzing our prompts, researchers can collect videos from YouTube according to users’ preferred subjects and directions. This avoids the previous problem, i.e., blindly expanding the training set led to wasted resources and low user satisfaction. (2) Improving evaluation practicality. Existing image-to-video benchmarks, such as $\mathtt{VBench}$-$\mathtt{I2V}$ [25], $\mathtt{I2V}$-$\mathtt{Bench}$ [49], and $\mathtt{AIGCBench}$ [18], suffer from a limited number of topics and prompts designed by experts, which may not accurately reflect real-world user needs. With the help of prompts in TIP-I2V, researchers can build a more comprehensive and practical benchmark for evaluating image-to-video models. For safer models: A major safety concern of image-to-video generation is misinformation, i.e., they can make objects or humans in images to perform actions they never did. For instance, given an image of $\mathtt{Elon}$ $\mathtt{Musk}$ and $\mathtt{Donald}$ $\mathtt{Trump}$ together, image-to-video models could generate a video of them fighting, misleading the public. To address this issue, videos in our TIP-I2V allows researchers to train a model to (1) distinguish between generated videos from images and real videos, and (2) trace the source image from any given frame in a generated video. Beyond these areas, we also encourage researchers to explore additional directions.
@@
Figure 2: A data point in our TIP-I2V includes UUID, anonymous UserID, text and image prompt, timestamp, subject and direction, NSFW status of text and image, text and image embedding, and the corresponding generated videos.
@@ 3 Curating TIP-I2V
Fig. 2 is a data point in the TIP-I2V. It includes a UUID, an anonymous UserID, text and image prompts, timestamp, subject and direction, NSFW status for text and images, embeddings for both text and images, and generated videos from five state-of-the-art image-to-video models. The following explains how we collect these pieces of information.
@@
Assigning UUIDs, UserIDs, timestamps, subjects, embeddings and NSFW status. To facilitate subsequent research utilizing our TIP-I2V, we (1) calculate UUIDs and anonymous UserIDs based on unique prompts and user names, respectively, (2) extract timestamps from Pika videos, (3) infer subjects and directions using $\mathtt{GPT}$-$\mathtt{4o}$ [42], (4) embed text and image prompts using $\mathtt{text}$-$\mathtt{embedding}$-$\mathtt{3}$-$\mathtt{large}$ [41] and $\mathtt{CLIP}$ [47], respectively, and (5) assign NSFW status to text and image prompts using $\mathtt{Detoxify}$ [54] and $\mathtt{nsfw\_image\_detection}$ [17], respectively.
@@ 5.5 Other Promising Research
$\bullet$ Cache for speeding-up generation. When image-to-video models are matured and consistently produce high-quality videos, researchers could use these models with our prompts to pre-generate and cache videos. Then, when a new prompt is received, the video can be generated quickly based on the cache, potentially saving several minutes.
@@ 7 Comparing TIP-I2V with Panda-70M
As shown in Fig. 8, we compare the text semantics of our TIP-I2V and Panda-70M [12] using WizMap [65] to highlight their differences.
@@ 14 Details of Fine-tuning VideoMAE
We fine-tune the Video Masked Autoencoder (VideoMAE) [53] on the TIP-ID dataset for a video classification task. Specifically, our preprocessing pipeline includes (1) temporal subsampling, (2) spatial transformations, and (3) normalization. During training, the spatial transformations consist of random short-side scaling, random cropping to $224\times 224$, and random horizontal flipping; for testing, we only resize frames to $224\times 224$. The pre-trained model is downloaded from the VideoMAE’s official Hugging Face repository, and we adjust it to match the number of classes in our dataset, i.e., 3, by updating the classification head. Training is distributed across a server with $8$ A100 GPUs. The mini-batch size is $8$, the learning rate is $5\times 10^{-5}$, and the number of iterations is $20,000$.
@@ 17 Potential Social Impact
The TIP-I2V dataset has potential for positive social impact by enhancing digital creativity and fostering responsible AI use. Specifically, by helping the creation of more user-responsive image-to-video models, TIP-I2V enables content creators to create engaging and customized videos. Additionally, TIP-I2V contributes to the development of detection models that help verify authenticity, trace image sources, and prevent harmful content misuse. Nevertheless, the TIP-I2V dataset may also have potential negative social impacts if misused. Below, we outline several potential negative social impacts and provide solutions:
@@
$\bullet$ Privacy. Although, per Pika’s regulations, users agree to make their input and output publicly available, some may still feel uncomfortable with the inclusion of their prompts in TIP-I2V. To enhance user privacy, we implement the following measures: (1) each prompt is assigned a new UUID and an anonymous UserID instead of the identifiable original user name; and (2) users have the right to request that their contributions be removed from TIP-I2V. They can simply email us to make this request.